# thanks: 0000-0002-5209-872Xthanks: 0000-0003-2028-6782

### v1

@@ Acknowledgements
We thank Christoph Ternes for helpful comments in the early stages of this project. PBD acknowledges support from the US Department of Energy under Grant Contract DE-SC0012704. This manuscript has been authored in part by FermiForward Discovery Group, LLC under Contract Number 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.
Appendix A Validation
Here we provide plots validating that the code behaves as expected. In fig. 9 we compare the result for day-time neutrino disappearance $\nu_{e}\to\nu_{e}$ probability produced in Sun at $\rho=100$ g/cc and $Y_{e}=\frac{2}{3}$ with the well known analytic approximation for two flavors [104] with some three-flavor information [105]:
with
Appendix B Determinant and Adjugate of $(\lambda I-H)$:
@@
For the adjugate of $(\lambda I-H)$ we need the real symmetric $S\equiv\text{Tr}[H]I-H$ and $T\equiv\text{Adj}[H]$ matrices, see [33],

### v2

@@ Acknowledgements
We thank Christoph Ternes for helpful comments in the early stages of this project, the anonymous referee for helpful comments, and Taku Izubuchi for a careful examination of the code. PBD acknowledges support from the US Department of Energy under Grant Contract DE-SC0012704. This manuscript has been authored in part by FermiForward Discovery Group, LLC under Contract Number 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.
Appendix A Determinant and Adjugate of $(\lambda I-H)$:
@@
For the adjugate of $(\lambda I-H)$ we need the real symmetric $S\equiv\text{Tr}[H]I-H$ and $T\equiv\text{Adj}[H]$ matrices [34],
@@
Appendix B Validation
Here we provide plots validating that the code behaves as expected.
In Figs. 10-11 we show the difference in probability between a detector on the surface and a detector 2 km deep for all four relevant oscillation channels in both the NO and the IO. The dominant features are horizontal trajectories at low energies. See also table 2 for a summary of these figures.
Figure 10: The difference in probabilities between a detector on the surface and a detector 2 km below the surface for disappearance (top) and appearance (bottom) probabilities across parameters typical for atmospheric oscillations in the normal ordering. The inverted ordering is shown in Fig. 11. The left (right) panels are for neutrinos (antineutrinos). The production height is 10 km. The largest effect is for low energies and near horizontal trajectories.
Figure 11: The same as Fig. 10 but in the IO.
In Fig. 12 we compare the result for day-time neutrino disappearance $\nu_{e}\to\nu_{e}$ probability produced in Sun at $\rho=100$ g/cc and $Y_{e}=\frac{2}{3}$ with the well known analytic approximation for two flavors [106] with some three-flavor information [107]:
with